# Akchen and Mišić

### v3

@@ 4.1 Benders reformulation of the SplitMIO relaxation
5:        Set $q_{s}\leftarrow x_{v(t,s)}-{\color[rgb]{0,0,0}\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{left}(s)}y_{t,\tau(j)}}$
@@
7:        Set $q_{s}\leftarrow 1-x_{v(t,s)}-{\color[rgb]{0,0,0}\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{right}(s)}y_{t,\tau(j)}}$
@@ 4.3.1 SplitMIO
$\alpha_{t,s}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{RS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$
$\beta_{t,s}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{LS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$
@@ 4.3.2 ProductMIO
$\alpha_{t,i}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{RP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$ (18)
$\beta_{t,i}$ =$=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{LP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right.$ (21)
@@ 4.4 Overall Benders algorithm
We conclude this section by highlighting the advantages of the SplitMIO formulation in the large-scale regime. At first glance, following Section 3, ProductMIO may appear preferable to SplitMIO since it is a tighter formulation (Proposition 3.5) with fewer constraints. However, in large-scale settings, the proposed primal-dual greedy algorithm enables significantly faster computation of the linear relaxation for SplitMIO compared to ProductMIO. As a result, as outlined above, SplitMIO can benefit from a more efficient warm start during the relaxation phase. Additionally, the objective value of the linear relaxation provides a valid upper bound for the assortment problem (2). This upper bound is particularly valuable in large-scale instances where solving SplitMIO or ProductMIO exactly may be impractical. A strong upper bound helps reduce the number of branches explored in the branch-and-bound procedure and serves as a benchmark for assessing the suboptimality of a given assortment solution. In Appendix 11.3, we demonstrate that, with the greedy algorithm, solving the relaxed SplitMIO problem can be up to 95% faster than solving the relaxed ProductMIO while providing comparable upper bounds.
@@ 5.1 Background
For all three types of forests, we generate the purchase decision $c(t,\ell)$ for each leaf $\ell$ in each tree $t$ in the following way: for each leaf $\ell$, we uniformly at random choose a product $i\in\cup_{s\in\mathbf{LS}(\ell)}\{v(t,s)\}\cup\{0\}$. In words, the purchase decision is chosen to be consistent with the products that are known to be in the assortment if leaf $\ell$ is reached. Figure 2 shows an example of each type of tree (T1, T2, and T3). Given a forest of any of the three types above, we generate the customer type probability vector $\boldsymbol{\lambda}=(\lambda_{t})_{t\in F}$ by drawing it uniformly from the $(|F|-1)-$dimensional unit simplex.
@@ 5.2 Experiment #1: Formulation Strength
We also test the tractability of SplitMIO and ProductMIO when they are solved as integer programs (i.e., not as relaxation). We present the results in Section 11.1.
@@ 5.3 Experiment #2: Benders Decomposition for Large-Scale Problems
In this experiment, we report on the performance of our Benders decomposition approach for solving large scale instances of SplitMIO. We focus on the SplitMIO formulation, as we are able to efficiently generate Benders cuts for both fractional and integral values of $\mathbf{x}$.
We deviate from our previous experiments by generating a collection of T3 instances with $n\in\{200,500,1000,2000,3000\}$, $|F|=500$ trees and $|\mathbf{leaves}(t)|=512$ leaves. As before, the marginal revenue $\rho_{i}$ of each product $i$ is chosen uniformly at random from $\{1,\dots,100\}$. For each value of $n$, we generate 5 instances. For each instance, we solve the SplitMIO problem subject to the constraint $\sum_{i=1}^{n}x_{i}=b$, where $b$ is set as $b=\mu n$ and we vary $\mu\in\{0.02,0.04,0.06,0.08,0.10,0.12\}$.
@@ 6.3 Case Study: Beer Category
We first observe that the consumer choice within the Beer category exhibits a notable phenomenon known as choice overload, a well-documented behavioral anomaly in the field of marketing science (Iyengar and Lepper 2000, Chernev et al. 2015, Long et al. 2023). This phenomenon describes the situation in which an abundance of available products can actually deter customers from making a purchase. Choice overload contradicts with the principles of rational choice theory. According to this theory, when a store increases the assortment size, a rational consumer should theoretically be more inclined to make a purchase from the assortment. This is because a larger assortment would increase the chance that the customer finds one product that aligns with her preference. In the context of choice modeling, a rational choice model satisfies the following:
@@
Note that equation (23) directly stems from the regularity property, which the ranking-based model and other RUM choice models, such as the mixed-MNL model, always adhere to (Rieskamp et al. 2006, Jagabathula and Rusmevichientong 2019). However, real-world scenarios often deviate from equation (23). When provided with more options, customers may become fatigued from the search process or experience a decrease in their confidence in decision-making, leading them to opt not to make a purchase. In the consumer psychology literature, Iyengar and Lepper (2000) conducted a field experiment demonstrating that individuals are more likely to purchase gourmet jams or chocolates when presented with a smaller assortment of size six rather than a more extensive assortment of size 24 or 30.
@@
Figure 3 (left) visually depicts the prevalence of choice overload within the Beer category of the IRI dataset. Each data point in the figure corresponds to a pair ($|S|$, $\bar{P}(0\mid S)$), where $|S|$ represents the assortment size of a historical assortment $S\in\mathcal{S}_{\text{hist}}\equiv\{S_{w,s}\mid w\in\mathcal{W},s\in\mathcal{S}\}$, and $\bar{P}(0\mid S)$ represents the empirical probability of the no-purchase option being chosen when the assortment $S$ was offered. For a better illustration, we have excluded assortments with very few transactions (less than or equal to 10) from Figure 3 (left). In this figure, the rightmost data point, denoted by a red triangle, corresponds to the no-purchase probability when the assortment contains all available products ($|S|=9$). It is important to note that in a scenario where equation (23) holds, indicating the absence of choice overload, each data point should exhibit a higher $y$-value than the red triangle. However, this condition is met for only 8 out of the 41 data points in the figure, suggesting a violation of equation (23) within the Beer category.
The presence of choice overload in the data has a significant impact on the estimation outcomes of both the ranking-based model and the decision forest model, leading to divergent assortment decisions. To begin, we define $\bar{P}_{\mathcal{C}}(0\mid m)$ as the average no-purchase probability when presented with an assortment of size $m$; formally, $\bar{P}_{\mathcal{C}}(0\mid m)={\sum_{S\subseteq\mathcal{N}\,:\,|S|=m}P_{\mathcal{C}}(0\mid S)}\big{/}{|\{S\subseteq\mathcal{N}\,:\,|S|=m\}|}$, where $P_{\mathcal{C}}(0\mid S)$ represents the no-purchase probability given assortment $S$ under a choice model $\mathcal{C}$. We consider two choice models, the decision forest model (DF) and the ranking-based model (RM), both estimated from data. Figure 3 (right) presents the average no-purchase probability curves, $\bar{P}_{\mathcal{C}}(0\mid m)$, for these two models, with the error bars representing the standard deviation. Recall that the ranking-based model adheres to the regularity property and strictly follows equation (23). Consequently, the resulting no-purchase probability curve (depicted in blue) does not exhibit the choice overload effect; it consistently decreases as we expand the assortment size. In contrast, the decision forest model is not bound by the regularity property and has the capacity to learn the choice overload effect from the data. The resulting no-purchase probability curve (shown in red) indicates that customers are indeed more inclined to make a purchase as the assortment initially expands. However, after reaching an assortment size of $|S|=6$, the choice overload effect becomes evident. Customers become less motivated, and we observe an increase in the no-purchase probability when $|S|\geq 7$.
@@ 7 Conclusions
We developed a mixed-integer optimization methodology for solving the assortment optimization problem under the decision forest model, which accounts for both rational and non-rational customer behaviors. We first established the inapproximability of the problem, then presented two formulations, linking them to the optimization models in the literature. To enhance scalability, we proposed a large-scale solution approach based on Benders decomposition. Through synthetic data experiments, we demonstrated the efficiency of our methods, and with real-world data, we showed how consumer behavioral anomalies can be learned and leveraged within our framework.
@@ Acknowledgments
We sincerely thank the department editor Huseyin Topaloglu, the associate editor and two anonymous reviewers for their thoughtful comments that have helped to significantly improve the paper.

### v4

@@ 4.1 Benders reformulation of the SplitMIO relaxation
5:    Set $q_{s}\leftarrow x_{v(t,s)}-\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{left}(s)}y_{t,\tau(j)}$
@@
7:    Set $q_{s}\leftarrow 1-x_{v(t,s)}-\sum_{1\leq j\leq i-1:\tau(j)\in\mathbf{right}(s)}y_{t,\tau(j)}$
@@ 4.3.1 SplitMIO
$\alpha_{t,s}$ $=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{RS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right$.
$\beta_{t,s}$ $=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(s)}r_{t,\ell}-r_{t,\ell^{*}}\}&\text{if}\ s\in\mathbf{LS}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right$.
@@ 4.3.2 ProductMIO
$\alpha_{t,i}$ $=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{left}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{RP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right$. (18)
$\beta_{t,i}$ $=\left\{\begin{array}[]{ll}\max\{0,\max_{\ell\in\mathbf{right}(i)}r_{t,\ell}-r_{t,\ell^{*}}\},&\text{if}\ i\in\mathbf{LP}(\ell^{*}),\\ 0&\text{otherwise},\end{array}\right$. (21)
@@ 4.4 Overall Benders algorithm
We conclude this section by highlighting the advantages of the SplitMIO formulation in the large-scale regime. At first glance, following Section 3, ProductMIO may appear preferable to SplitMIO since it is a tighter formulation (Proposition 3.4) and involves no more constraints. However, in large-scale settings, the proposed primal-dual greedy algorithm enables significantly faster computation of the linear relaxation for SplitMIO compared to ProductMIO. As a result, as outlined above, SplitMIO can benefit from a more efficient warm start during the relaxation phase.
To support the discussion surrounding the two-phase structure and the advantages afforded by SplitMIO through the greedy algorithm, we provide two pieces of numerical evidence. First, Appendix 12.4 shows that incorporating the relaxation phase can effectively reduce the optimality gap of the Benders approach, especially for large instances. Second, Appendix 12.5 demonstrates that, with the greedy algorithm, solving the relaxed SplitMIO problem can be up to 95% faster than solving the relaxed ProductMIO while providing comparable upper bounds.
@@ 5.1 Background
For all three types of forests, we generate the purchase decision $c(t,\ell)$ for each leaf $\ell$ in each tree $t$ in the following way: for each leaf $\ell$, we uniformly at random choose a product $i\in\cup_{s\in\mathbf{LS}(\ell)}\{v(t,s)\}\cup\{0\}$. In words, the purchase decision is chosen to be consistent with the products that are known to be in the assortment if leaf $\ell$ is reached. Figure 2 shows an example of each type of tree (T1, T2, and T3). Given a forest of any of the three types above, we generate the customer type probability vector $\boldsymbol{\lambda}=(\lambda_{t})_{t\in F}$ by drawing it uniformly from the (|F|1)(|F|-1)-dimensional unit simplex.
@@ 5.2 Experiment #1: Formulation Strength
We also test the tractability of SplitMIO and ProductMIO when they are solved as integer programs (i.e., not as a relaxation). We present the results in Appendix 12.2.
We note that the experiments in this section (Experiment #1) and in Appendix 12.2 focus on instances where the full formulations of SplitMIO and ProductMIO fit in memory and can be solved directly by Gurobi. In this setting, both Table 1 and Table 7 in Appendix 12.2 show that ProductMIO is generally preferable to SplitMIO. In particular, when trees are highly structured with extreme product repetition in split nodes, as in the T1 instances, ProductMIO exhibits a much smaller integrality gap and, in some cases, is solved to optimality by Gurobi up to five times faster than SplitMIO. When the trees are less repetitive, as in the T2 and T3 instances, the two formulations perform comparably in terms of both integrality gap and runtime. In the next experiment, we further scale up the problem size to enter a regime where directly solving the full formulations of SplitMIO and ProductMIO with Gurobi is no longer feasible.
@@ 5.3 Experiment #2: Benders Decomposition for Large-Scale Problems
In this experiment, we report on the performance of our Benders decomposition approach for solving large scale instances of SplitMIO, where solving the full extended formulation may no longer be viable. These problem instances are significantly larger than those considered in the previous experiment. We focus on the SplitMIO formulation, as we are able to efficiently generate Benders cuts for both fractional and integral values of $\mathbf{x}$. Two further experiments in this large-scale regime are presented in the Appendix. Appendix 12.4 highlights the value of the Benders cuts generated from fractional values of $\mathbf{x}$ within the full Benders decomposition framework. Appendix 12.5 compares SplitMIO and ProductMIO, demonstrating that, with the proposed primal–dual greedy algorithm in Section 4.1, the linear relaxation of SplitMIO can be solved up to 95% faster than that of ProductMIO.
First, we scale up the problem sizes considered in Experiment #1. Specifically, we generate a collection of T3 instances with $n\in\{200,500,1000,2000,3000\}$, $|F|=500$ trees and $|\mathbf{leaves}(t)|=512$ leaves. As before, the marginal revenue $\rho_{i}$ of each product $i$ is chosen uniformly at random from $\{1,\dots,100\}$. For each value of $n$, we generate 5 instances. For each instance, we solve the SplitMIO problem subject to the constraint $\sum_{i=1}^{n}x_{i}=b$, where $b$ is set as $b=\mu n$ and we vary $\mu\in\{0.02,0.04,0.06,0.08,0.10,0.12\}$.
@@ 6.3 Case Study: Beer Category
We first observe that the consumer choice within the Beer category exhibits a notable phenomenon known as choice overload, a well-documented behavioral anomaly in the field of marketing science (Iyengar and Lepper 2000, Chernev et al. 2015, Long et al. 2023). This phenomenon describes the situation in which an abundance of available products can actually deter customers from making a purchase. Choice overload contradicts the principles of rational choice theory. According to this theory, when a store increases the assortment size, a rational consumer should theoretically be more inclined to make a purchase from the assortment. This is because a larger assortment would increase the chance that the customer finds one product that aligns with her preference. In the context of choice modeling, a rational choice model satisfies the following:
@@
Note that Condition (23) directly stems from the regularity property, which the ranking-based model and other RUM choice models, such as the mixed-MNL model, always adhere to (Rieskamp et al. 2006, Jagabathula and Rusmevichientong 2019). However, real-world scenarios often deviate from Condition (23). When provided with more options, customers may become fatigued from the search process or experience a decrease in their confidence in decision-making, leading them to opt not to make a purchase. In the consumer psychology literature, Iyengar and Lepper (2000) conducted a field experiment demonstrating that individuals are more likely to purchase gourmet jams or chocolates when presented with a smaller assortment of size six rather than a more extensive assortment of size 24 or 30.
@@
Figure 3 (left) visually depicts the prevalence of choice overload within the Beer category of the IRI dataset. Each data point in the figure corresponds to a pair ($|S|$, $\bar{P}(0\mid S)$), where $|S|$ represents the assortment size of a historical assortment $S\in\mathcal{S}_{\text{hist}}\equiv\{S_{w,s}\mid w\in\mathcal{W},s\in\mathcal{S}\}$, and $\bar{P}(0\mid S)$ represents the empirical probability of the no-purchase option being chosen when the assortment $S$ was offered. For a better illustration, we have excluded assortments with very few transactions (less than or equal to 10) from Figure 3 (left). In this figure, the rightmost data point, denoted by a red triangle, corresponds to the no-purchase probability when the assortment contains all available products ($|S|=9$). It is important to note that in a scenario where Condition (23) holds, indicating the absence of choice overload, each data point should exhibit a higher $y$-value than the red triangle. However, this condition is met for only 8 out of the 41 data points in the figure, suggesting a violation of Condition (23) within the Beer category.
The presence of choice overload in the data has a significant impact on the estimation outcomes of both the ranking-based model and the decision forest model, leading to divergent assortment decisions. To begin, we define $\bar{P}_{\mathcal{C}}(0\mid m)$ as the average no-purchase probability when presented with an assortment of size $m$; formally, $\bar{P}_{\mathcal{C}}(0\mid m)={\sum_{S\subseteq\mathcal{N}\,:\,|S|=m}P_{\mathcal{C}}(0\mid S)}\big/\penalty 50{|\{S\subseteq\mathcal{N}\,:\,|S|=m\}|}$, where $P_{\mathcal{C}}(0\mid S)$ represents the no-purchase probability given assortment $S$ under a choice model $\mathcal{C}$. We consider two choice models, the decision forest model (DF) and the ranking-based model (RM), both estimated from data. Figure 3 (right) presents the average no-purchase probability curves, $\bar{P}_{\mathcal{C}}(0\mid m)$, for these two models, with the error bars representing the standard deviation. Recall that the ranking-based model adheres to the regularity property and strictly follows Condition (23). Consequently, the resulting no-purchase probability curve (depicted in blue) does not exhibit the choice overload effect; it consistently decreases as we expand the assortment size. In contrast, the decision forest model is not bound by the regularity property and has the capacity to learn the choice overload effect from the data. The resulting no-purchase probability curve (shown in red) indicates that customers are indeed more inclined to make a purchase as the assortment initially expands. However, after reaching an assortment size of $|S|=6$, the choice overload effect becomes evident. Customers become less motivated, and we observe an increase in the no-purchase probability when $|S|\geq 7$.
@@
6.4 Additional Numerical Experiments based on the Sushi Dataset
The IRI dataset provides valuable real-world transaction data, but it also comes with limitations. First, it does not specify a ground-truth choice model, so the evaluation of assortment performance must rely on comparisons across models, as in Table 3. Second, even after standard preprocessing steps from the literature, the historical assortments and no-purchase options are not always fully consistent across all assortment–choice pairs. To address these limitations, we complement the IRI-based analysis with a semi-synthetic experiment using the Sushi dataset (Kamishima 2003), presented in Section 13.
In this experiment, we construct a ground-truth choice model over ten sushi types ($n=10$) that combines (i) the ranking-based preferences in the Sushi dataset and (ii) a synthetic choice overload effect, where customers increasingly opt for no purchase as assortments expand. We use this ground truth to generate synthetic assortment–choice data, from which we then estimate the MNL, ranking-based, and decision forest models. Let $S^{\text{MNL}}$, $S^{\text{RB}}$, and $S^{\text{DF}}$ denote the optimal assortments recommended by the respective estimated models. We then evaluate the expected revenues of these assortments under the synthetic ground truth.
The results show that the decision forest model ($S^{\text{DF}}$) consistently delivers strong revenue performance, even when the choice overload effect is intensified. By contrast, the revenues of $S^{\text{MNL}}$ and $S^{\text{RB}}$ degrade significantly under stronger overload. When the overload effect is mild, the ranking-based model performs best—as expected, since the ground truth is nearly ranking-based in this case. However, as overload increases, the decision forest model significantly outperforms both benchmarks.
This semi-synthetic experiment highlights the end-to-end (data-to-decision) performance of the decision forest model in the presence of a specific behavioral anomaly. At the same time, it isolates only one anomaly—choice overload—whereas real-world settings often exhibit multiple behavioral patterns simultaneously, such as the decoy effect (Huber et al. 1982) and compromise effect (Simonson 1989). In such cases, the IRI dataset study in Section 6.2 provides a broader perspective on how the decision forest model compares with rational-choice models such as the ranking-based model. Taken together, the IRI and Sushi experiments complement each other: the former grounds our analysis in real data with rich behavioral heterogeneity, while the latter offers a controlled environment with a known ground truth. Together, they provide a comprehensive assessment of the decision forest model.
@@ 7 Conclusions
We developed a mixed-integer optimization methodology for solving the assortment optimization problem under the decision forest model, which accounts for both rational and non-rational customer behaviors. We first established the inapproximability of the problem and then presented two mixed-integer programs. To enhance scalability, we proposed a large-scale solution approach based on Benders decomposition. Through synthetic data experiments, we demonstrated the efficiency of our methods, and with real-world data, we showed how consumer behavioral anomalies can be learned and leveraged within our framework.
@@ Acknowledgments
We sincerely thank the department editor Huseyin Topaloglu, the associate editor and two anonymous reviewers for their thoughtful comments that have helped to significantly improve the paper. This work has benefited from the generous research support provided by the UCL School of Management and the UCLA Anderson School of Management.